# Comparison of Generative Learning Methods for Turbulence Modeling

### v1

@@ 3.3.2 Continuous Limit
for some ${\epsilon_{t}}$, each independently identically distributed by the unit normal. It is natural to ask what happens in the continuous limit, namely as $T$ becomes very large. [87, 89, 90] show that we arrive at the stochastic differential equation (SDE)
$dX_{t}=-\frac{1}{2}\beta(t)X_{t}dt+\sqrt{\beta(t)}dW_{t}$ (17)
Where $\beta(t)$ is an appropriate continuation of the discrete noise schedule, and $W_{t}$ is the standard Wiener process. The reverse is also true, as discretising this equation again returns us to the case above [90]. In this sense SDEs can be seen as an overarching framework for diffusion models. Given that the noise schedule increases sufficiently quickly so that $x_{T}\approx\mathcal{N}(0,I)$, the time reversed (see the main theorem in [91]) SDE is then
this is known as a ‘variance preserving’ (VP) SDE. The probability distribution $p(\cdot;x_{t})$ (here $t$ is a continuous parameter) satisfies the Fokker-Planck equation, also known as Kolmogorov’s forwards equation, which is an ODE describing how the probability density decays with time [87]. This equation, like equation 18, can be used to sample from the original data distribution [89].
3.3.3 Score-Matching
Another possible interpretation of Diffusion models is to understand them by analogy with score-matching methods, which learn the ‘score’ (the gradient of log probability) with methods such as Langevin dynamics [92]. Note that the score is known for a Gaussian. Indeed, in the formalism we have used so far,
$\nabla\ln{\mathcal{N}(x;\,\mu,\sigma^{2})}=\frac{\mu-x}{\sigma^{2}}$ (19)
Apply the reparameterisation trick from 3.1 to the distribution of $x_{t}$ and obtain
Learning this score function with mean-squared error leads to the same optimisation problem as 15. Philosophically, this is much the same as models formulated under the VAE interpretation. Where the schools of thought differ is in the noising process (which in turns leads to a different sampling process). SBM (score-based matching) models utilise
[…]
for some noise scales $\{\sigma_{t}\}$ typically set as a geometric sequence, which again ensure in the limit that the fully noised distribution is equivalent to $\mathcal{N}(0,I).$ Performing the limit procedure above on this process also leads to a SDE, known as a ‘variance exploding’ (VE) SDE, a careful treatment of which can be found in [87].
4 Dataset
4.1 Setup of simulation
We perform our experiments on the test case of a flow around a cylinder at Reynolds number 3900 which is well studied in the literature [93, 94, 95, 93, 96, 97]. The flow field is characterized by a Kármán vortex street developing in the wake of the cylinder. The vortex street consists of a characteristic coherent vortex system in which the rotational axes of the individual vortices are aligned with the cylinder axis.
The data set of grayscale images (see figure 4) was generated by post-processing the transient LES velocity field data using a projection mapping in the sense that the system remains ergodic on a reduced state space. Let $V(\xi,t)=(V_{x}(\xi,t),V_{y}(\xi,t),V_{z}(\xi,t))$ be the velocity field of the fluid. Then for our dataset, the gray scale shows the distribution of the absolute deviation of the local fluctuating velocity magnitude $c(\xi,t)=\sqrt{V_{x}(\xi,t)^{2}+V_{y}(\xi,t)^{2}+V_{z}(\xi,t)^{2}}$ at the location $\xi$ from its time average $c^{\prime}(\xi,t)=|c(\xi,t)-\overline{c}(\xi)|$ with ${}~{}\overline{c}(\xi)=\frac{1}{T}\int_{0}^{T}c(\xi,t)\,\mathrm{d}t$. For the numerical setup of the LES, see [28]. In total, the data set consists of $100,000$ images with a resolution of $1000\times 600$ pixels. The full data set is available at [98].
@@ 5.1 Implementation details
In this work, we investigate VAE and DDPM as alternative generative learning approaches for turbulence modeling and compare them to the DCGAN turbulent flow generator developed in our previous work [28], where the implementation details can also be found. For the training of the VAE and the DDPM, the architecture proposed by [23] and [25], respectively, was adopted. In particular, the VAE consists of an encoder and decoder of similar structure with seven hidden layers containing $\Sigma_{n=3}^{9}{2^{n}}$ neurons. We used a more advanced DDPM model with attention, which means it is combined with a transformer [99] and shadowed by an Exponential Moving Average (EMA) model [100]. Both, VAE and DDPM take images of size $m\times m,m\in\mathbb{N}^{+}$ as input. Therefore, the images were pre-processed prior to training by removing excess white pixels in front of the static wake (150 pixels in the width directory) and by compressing them to a resolution of $512\times 512$. Finally, all generative networks were built and trained with PyTorch [101]. A summary and comparison of the relevant training parameters, such as learning rate or effective batch size, of the three generative models investigated can be found in section 5.1.
[…]
=1.2mm {tabu}|l|l|l|l| Model VAE DCGAN DDPM (with attention)
@@
Optimiser Adam Adam Adam
@@
Training Images $100,000$ $5,000$ $100,000$
Image Resolution $512\times 512$ $512\times 512$ $512\times 512$
Table 1: Summary of the parameter settings of the generative models VAE, DCGAN and DDPM that were investigated. Note that the effective batch size is the actual batch size multiplied by the gradient accumulation.
@@ 5.2 Physics-based evaluation metrics
Figure 5: Areas considered for physics-based evaluation. The vertical blue line shows the starting point directly behind the vertebral body. The first reddish vertical line represents the reference line for the small evaluation grid of 12 pixels in $x$ direction away from the starting point (pixel value 300 in the image). The second reddish vertical line describes the reference line for the wide evaluation grid, which inherits 25% of the turbulent flow behind the cylinder (pixel value 500 in the image).
To numerically compare the results obtained by the considered generative models, we compute two metrics - the mean pixel values and the variance of the local fluctuating velocity magnitude $c(\xi,t)$ at point $\xi$ in two areas immediately behind the cylinder, where the strongest statistical fluctuation of the velocities is given (see figure 5). For a detailed description and justification of why these metrics are appropriate for the test case of the flow around a cylinder, see our preliminary work [28]. We performed the evaluations with $5,000$ images of the LES and $5,000$ images synthesized by the VAE, DCGAN, and DDPM, respectively.
It is worth noting that no known method is capable of reproducing exactly the turbulent flow for a given set of initial conditions, if such a thing even exists, and instead, produce stochastic realisations of what the flow may look like. All of these realisations contain some sort of error, and as such, the models trained on this flawed data will never be able to improve on the error that the data contains. Thus, we do not claim to be able to approximate the flow itself, but rather, we are attempting to approximate the LES data.
@@ 5.3 Results Discussion
Figure 6 shows example outputs generated by the three different models, highlighting the variance in quality. The VAE model is clearly the weakest performer, as its generated images are visibly of lower quality compared to those of the other models. It places the wake in the correct position and synthesizes some rough structure of the turbulent flow, but the images are darker, blurred, and lack detail in the vortices. However, when comparing the LES dataset, DCGAN, and DDPM outputs, the differences are not immediately apparent to the naked eye.
Instead, we need to turn to figure 7 and figure 8, which show the physics-based metrics measured along the $y$-axis over a grid in the $x$-direction, ending at the 300px and 500px mark, respectively (see figure 5). $\delta$ corresponds to the bandwidth of the vortex body, so that $y/\delta$ defines the normalized pixel index value of the $y$-axis. These figures confirm the initial visual assessment, namely, the VAE significantly underperforms, as evidenced by its overprediction of mean pixel values in figure 6(a) and figure 7(a) and underprediction of the variation of the pixel values in figure 6(b) and figure 7(b). Clearly, the VAE struggles to accurately represent the statistical properties of the flow.
A noticeable issue with the DCGAN output is the presence of a darkened background in a non-trivial proportion of the images, represented by slightly offset tails in the mean pixel value distributions. The background should be uniformly white (with pixel values exactly zero). This deviation is likely due to the DCGAN’s tendency to prioritize the generation of distinguishable prominent features, sometimes at the expense of accurately capturing uniform background regions, since the discriminator may learn to prioritize these high entropy regions as better predictors of whether or not the image is artificially generated. On top of this, the model inherently introduces noise as part of the generation process. Since the background should be identically zero, the noise can manifest much more noticeably in this region. The DDPM images also suffer from this problem for the same reason, but to a much lesser extent.
Both the GAN and DDPM models exhibit a much better alignment with the dataset with respect to this metric. Despite this, neither model is flawless. The DCGAN somewhat outperforms the DDPM, particularly in capturing the chaotic nature of the flow, since the DDPM tends to underpredict the variation in pixel values, indicating that it has difficulty fully capturing the chaotic turbulence. While the DCGAN also exhibits this underprediction, it does so to a lesser extent. In particular, the GAN better captures the shape of the distributions. An examination of the central region, found in figure 7(b), shows that the DCGAN is better able to recreate the “kinks” of the distribution, and in the majority of cases performs better than the DDPM at hitting the peaks.
(a) Mean pixel values
(b) Statistical fluctuation of deviation from mean pixel values
Figure 7: Comparison of (a) the mean pixel values and (b) the statistical fluctuation of the mean pixel values, over the grid ending at the 300px mark (see figure 5). Here, a pixel value represents the difference between the velocity at that point and the background. All datasets were normalized before evaluation.
(a) Mean pixel value
(b) Statistical fluctuation of deviation from mean pixel values
Figure 8: Comparison of (a) the mean pixel values and (b) the statistical fluctuation of the mean pixel values with a zoomed view, over the grid ending at the 500px mark (see figure 5). Here, a pixel value represents the difference between the velocity at that point and the background. All datasets were normalized before evaluation.
@@ 5.4 Computational Cost
We consider the training and inference time to measure the computational cost (see section 5.4). Since each of the considered generative models has its own properties, summarized in section 5.1, each model was trained on different amounts of A100 GPUs with 80GB each.
While VAE and DDPM needed all $100,000$ images to train, DCGAN training can be done with $0.05\%$ of the whole dataset without losing generalizability. In addition, the DDPM has over 135 million model parameters. To meet the challenge of a large dataset of high-resolution images and large models, VAE and DDPM were parallelized across multiple GPUs to minimize training time. Note that the speedup for training with more GPUs in parallel is almost, but slightly less than, linear. Nevertheless, the DCGAN was the fastest to train, despite being the model with the most parameters and being trained over 10 times more epochs than the VAE and DDPM on only one GPU. In terms of inference time, the VAE is significantly faster than the DCGAN and the DDPM, but the generated results don’t reach the quality of the LES. Therefore, the inference times of DCGAN and DDPM are of particular interest to us. Here we can see that DCGAN is also significantly faster than DDPM with 0.001 sec/frame compared to 36.3 sec/frame that DDPM needs to generate an image.
[…]
=1.2mm {tabu}|l|l|l|l| Model VAE DCGAN DDPM (with attention)
@@
Figure 9: A Kiviat diagram comparing the three generative models. Inference time and memory use are measured per sample, and the errors are measured by taking a mean square error between the respective error curves, and are then placed on a log-scale.

### v2

@@ 4.2 Computational Cost
At this stage, it is worth briefly addressing the computational effort involved in the datasets presented in this paper, as this is a key factor determining the feasibility of high-resolution LES and experimental PIV measurements.
4.2.1 Flow around a cylinder
@@
4.2.2 Flow behind array of seven cylinders
With the above-mentioned experiment, the data were collected at a 200 Hz frame rate, which converts to a measurement time of 15 s. However, the subsequent image processing took several hours on a desktop PC equipped with an Intel(R) Xeon(R) W-3223 CPU @ 3.50 GHz. In general, the advantage of experiments is that the processing time is not dependent on the Reynolds number, and significantly larger Reynolds numbers (i.e., higher levels of turbulence) can be characterized in a shorter time as compared to simulations. Major difficulties are the precise control of the boundary conditions and the fact that any measurement comes with an uncertainty that the AI methods need to deal with.
@@ 5.1 Implementation Details
Model VAE DCGAN DDPM (with attention)
@@
Optimizer Adam Adam Adam
@@
Training Images
100,000 (A)
   3,000 (B)
5,000 (A)
3,000 (B)
100,000 (A)
   3,000 (B)
Image Resolution
512×512 (A)
256×256 (B)
512×512 (A)
128×128 (B)
512×512 (A)
256×256 (B)
Table 1: Summary of the parameter settings of the generative models VAE, DCGAN, and DDPM, investigated using two datasets: (A) flow around a cylinder and (B) flow behind array of seven cylinders. The effective batch size is the actual batch size multiplied by the gradient accumulation.
In this work, we investigate VAE and DDPM as alternative generative learning approaches for turbulence modeling and compare them to the DCGAN turbulent flow generator developed in our previous work [drygala2022generative], where the implementation details can also be found. For the training of the VAE and the DDPM, the architecture proposed by [kingma2013] and [ho2020denoising], respectively, was adopted. In particular, the VAE consists of an encoder and decoder of similar structure with seven hidden layers containing $\Sigma_{n=3}^{9}{2^{n}}$ neurons. We used a more advanced DDPM model with attention, which means it is combined with a transformer [vaswani2023attentionneed] and shadowed by an Exponential Moving Average (EMA) model [Karras2023AnalyzingAI]. The generative models require square images as input for training. The images were resized to a resolution of $m\times m,m\in\mathbb{N}^{+}$ pixels. For the flow-around-a-cylinder dataset, an additional pre-processing step removed 150 pixels in front of the static wake where the flow is laminar and no turbulence is present. This resizing did not alter the LES- or PIV-resolved scales and represented a purely computational adjustment to meet the input requirements of the neural networks. Only the first $5,000$ images of the cylinder dataset were considered for GAN training. Finally, all generative networks were built and trained with PyTorch [pytorch2019]. A summary and comparison of the relevant training parameters, such as learning rate or effective batch size, of the three generative models investigated can be found in table˜1. At inference time, the input for all three generative models is a latent vector drawn from a simple prior distribution - here, the standard Gaussian - with the same dimensionality as during training (see table˜1). This latent vector is mapped by the trained generator to a turbulent flow field sample. In GANs and VAEs through a single forward pass, and in DDPMs through an iterative denoising process. Because training was performed without a temporal component, the generative models do not capture time dependencies but instead produce temporally independent state snapshots of turbulent flows.
@@ 5.2 Physics-based Evaluation Metrics
(a) Flow around a cylinder
(b) Flow behind array of seven cylinders.
Figure 6: Evaluation regions for both datasets, where the vertical blue line marks the start of the investigated wake behind the cylinders. For the flow around a single cylinder (a), the red lines indicate regions located at $6D$ (small near-wake region), $9D$, $13D$, $16D$, and $20D$, corresponding to approximately 25%, 50%, 75%, and 100% of the turbulent wake. For the flow behind an array of seven cylinders (b), these regions are located at $12.6D$, $25.6D$, $37.5D$, and $50D$. The specific locations within the numerical domains are computed separately for each dataset.
To assess the performance of the generative models, we compute two physically interpretable metrics - the mean local velocity fluctuation magnitude and its variance - at each spatial location $\xi$, across the turbulent wake.
For both datasets, the LES flow around a single cylinder and the PIV flow behind an array of seven cylinders, the evaluation is performed by progressively increasing the portion of the downstream wake considered. Specifically, we analyze 25%, 50%, 75%, and 100% of the turbulent wake, allowing us to assess model performance across regions of varying turbulence intensity (see figure˜6. In addition, for the LES dataset, we include a very small near-wake region located at $6D$ within the numerical domain, directly behind the cylinder, where turbulence intensity is highest and coherent vortex shedding dominates the flow dynamics.
For the PIV dataset, the velocity components in the streamwise and transverse directions ($u$ and $v$) are combined into the local velocity magnitude, $\sqrt{u^{2}+v^{2}}$, to provide a scalar measure of instantaneous velocity fluctuations, analogous to the quantity computed for the LES data.
The evaluation is based on 5,000 LES snapshots and 5,000 images generated by each of the VAE, DCGAN, and DDPM models for the single-cylinder case, and on 3,000 PIV snapshots and 3,000 synthesized images for each test case of the cylinder-array flow. These selections ensure that both near- and far-wake turbulence are captured, enabling a consistent and physically meaningful comparison between the model-generated and reference flows across increasing wake regions.
A detailed discussion of the physical relevance and justification for these metrics in cylinder wake flows is provided in [drygala2022generative].
It is worth noting that no known method is capable of reproducing exactly the turbulent flow for a given set of initial conditions, if such a thing even exists, and instead, produce stochastic realizations of what the flow may look like. All of these realizations contain some sort of error, and as such, the models trained on this flawed data will never be able to improve on the error that the data contains. Thus, we do not claim to be able to approximate the flow itself, but rather, we are attempting to approximate the LES and PIV data.
@@ 5.3 Results Discussion
Below, we discuss the results of the conducted experiments. For improved readability, only the evaluation results of the near-wake regions are included in the main part of the discussion. Additional evaluation plots are provided in appendix˜A.
5.3.1 Flow around a cylinder
@@
Figure˜7 shows example outputs generated by the three models, highlighting differences in their ability to reproduce turbulent flow structures. The VAE model is the weakest performer, as it reproduces the large-scale wake location but underestimates the amplitude of local velocity fluctuations and underrepresents small-scale vortical structures.
Differences between the LES, DCGAN, and DDPM outputs are less apparent by visual inspection alone. Therefore, we turn to the physics-based evaluation metrics shown in figure˜8 and figure˜9, which display the mean local velocity fluctuation magnitude and its variance along the $y$-axis across the downstream wake, evaluated at a small region directly behind the cylinder ending at position $6D$ in the numerical domain, and a near-wake region capturing up to 25% of the downstream wake ending at position $9D$. These figures confirm the earlier assessment: the VAE systematically overestimates the mean local fluctuation magnitude while underpredicting the variance, indicating that it fails to capture both the average turbulent intensity and the statistical fluctuations associated with coherent and small-scale turbulence.
(a) Mean local velocity fluctuations
(b) Variance of local velocity fluctuations
Figure 8: Comparison of (a) the mean local velocity fluctuation magnitude and (b) the variance of the local velocity fluctuations, evaluated over the near-wake region extending to $6D$ within the numerical domain behind the cylinder (see figure˜6(a)). Each value represents the deviation of the local flow velocity from the background or mean flow. All datasets were normalized prior to evaluation.
(a) Mean local velocity fluctuations
(b) Variance of local velocity fluctuations
Figure 9: Comparison of (a) the mean local velocity fluctuation magnitude and (b) the variance of the local velocity fluctuations, with a zoomed view, evaluated over the wake region comprising 25% of the wake flow directly behind the cylinder, extending to $9D$ within the numerical domain (see figure˜6(a)). Each value represents the deviation of the local flow velocity from the background or mean flow. All datasets were normalized prior to evaluation.
A noticeable issue with the DCGAN outputs is the presence of artificially elevated velocity fluctuations in regions of the flow that should remain laminar, corresponding to areas outside the primary wake structures. These deviations are manifested as slight offsets in the tails of the mean fluctuation distributions. The background flow in these regions is physically nearly steady, so any nonzero fluctuations reflect artifacts introduced by the generative model rather than actual turbulence. This behavior likely arises from the DCGAN’s emphasis on generating prominent vortical structures, sometimes at the expense of accurately representing uniform, laminar regions, since the discriminator may prioritize regions with more variability as indicators of realism. Additionally, the generative process inherently introduces small-scale noise, which becomes more noticeable in these near-laminar areas. The DDPM outputs show a similar effect, but the magnitude of the artificial fluctuations in the laminar background is smaller.
Both DCGAN and DDPM better reproduce the mean and variance of the LES flow, though neither is perfect. The DCGAN captures the chaotic nature of the turbulence more effectively, particularly in the near wake, as seen in its ability to reproduce peaks and inflection points in the variance distributions (figure˜9(b)). The DDPM underpredicts variance for high-intensity turbulent regions, indicating a reduced representation of extreme turbulent events.
Considering the near-wake regions, the DCGAN slightly outperforms the DDPM in reproducing areas of high-intensity velocity fluctuations. Examining the downstream evolution of the variance, particularly in the far wake beyond $16D$ (see figure˜16 and figure˜17), both models tend to underestimate the highest-intensity fluctuations, with this underestimation becoming most pronounced in the final 25% of the far-wake region. Interestingly, while the DCGAN captures the overall shape of the LES variance curve, it underestimates high-intensity turbulence in the far wake more than the DDPM. This suggests that the DCGAN better reproduces near-wake dynamics, whereas the DDPM retains somewhat higher fidelity in representing intense turbulent fluctuations further downstream.
5.3.2 Flow behind array of seven cylinders
Figure 10: Examples from the 4V2H test case showing, from left to right, the fully trained VAE, DCGAN, and DDPM, with the PIV dataset on the right for comparison.
(a) Mean local velocity fluctuations
(b) Variance of local velocity fluctuations
Figure 11: Comparison of (a) the mean local velocity fluctuation magnitude and (b) the variance of the local velocity fluctuations, evaluated over the wake region comprising 25% of the wake flow behind the array of seven cylinders in the 4V2H test case, extending to $12.6D$ within the numerical domain (see figure˜6(b)). Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
For the flow behind the array of seven cylinders, two test cases, 4V2H and 2V8H, were investigated. Starting with the 4V2H configuration, the results reveal behavior consistent with observations from LES data. In figure˜10, a visual inspection of the streamwise velocity component $u$ shows that the VAE captures the large-scale structures of the downstream wake but struggles to resolve finer vortex details. In contrast, both the DCGAN and the DDPM are able to reproduce small-scale vortex features more accurately. It can be observed, however, that the DCGAN-generated wakes contain noticeable noise, which can be traced back to the inherent measurement noise in the experimental PIV data. During training, the DCGAN appears to have learned that these small-scale fluctuations are characteristic of realistic samples and therefore reproduces them in its generated outputs.
The evaluation metrics support these qualitative impressions. Considering the near-wake region (see figure˜11), comprising the first 25% of the total wake length, the VAE tends to locally underestimate the mean local velocity fluctuation magnitude and fails to fully capture the statistical variability associated with coherent and small-scale turbulence. For the 4V2H case, both DCGAN and DDPM closely match the statistical properties of the measured PIV data. The DCGAN performs particularly well in the lower wake region, where the flow exhibits a large, steady vortex with weak fluctuations. Nevertheless, in the second high-intensity turbulence region, the DDPM slightly outperforms the DCGAN. Interestingly, in the far-wake region, comprising the final 25% of the wake, all generative surrogates tend to underestimate again the highest-intensity fluctuations, with this underestimation becoming most pronounced toward the end of the wake (see figure˜19).
A similar trend is observed for the second test case, 2V8H, confirming the general validity of these findings (see figure˜12). As described in figure˜13, the VAE captures the mean local fluctuation magnitude more accurately for the first time, while the DCGAN again outperforms the DDPM in reproducing the peak fluctuation levels. Since the measured PIV data of this case exhibit weaker high-intensity turbulence in the far-wake region, the quality of the generated data remains high across all evaluated regions (see figure˜20 and figure˜21).
@@ 5.4 Computational Cost
We consider the training and inference time to measure the computational cost (see table˜2). Since each of the considered generative models has its own properties, summarized in table˜1, each model was trained on different amounts of A100 GPUs with 80GB each.
While VAE and DDPM needed all $100,000$ images to train, DCGAN training can be done with $0.05\%$ of the whole dataset without losing generalizability. In addition, the DDPM has over 135 million model parameters. To meet the challenge of a large dataset of high-resolution images and large models, VAE and DDPM were parallelized across multiple GPUs to minimize training time. Note that the speedup for training with more GPUs in parallel is almost, but slightly less than, linear. The DCGAN was the fastest model to train, despite having the most parameters and being trained for more than ten times as many epochs as the VAE and DDPM, but on a single GPU. In terms of inference time, the VAE is significantly faster than the DCGAN and the DDPM, but the generated results don’t reach the quality of the LES. Consequently, the inference performance of DCGAN and DDPM is of particular interest. The DCGAN generates one snapshot in $0.001\,\text{s}$, whereas the DDPM requires $36.3\,\text{s}$ per snapshot. Each generated sample represents a state snapshot of the flow field drawn from the distribution of the LES dataset, which was originally sampled at $68.9\,\text{Hz}$ (i.e., every $\approx 250$ LES time steps, $\Delta t=5.8\times 10^{-5}\,\text{s}$). Thus, the reported “seconds per frame” values purely represent wall-clock inference times and are not related to physical time scales.
Model VAE DCGAN DDPM (with attention)
@@
For the experimental PIV data, computational cost plays a minor role, and we therefore focus our analysis on the LES cases. Nevertheless, the Kiviat diagram (see figure˜15) confirms that the considered generative models remain computationally advantageous also for the PIV data, while the main challenge in this case lies in handling the measurement uncertainties inherent to experimental observations.
Figure 12: Examples from the 2V8H test case showing, from left to right, the fully trained VAE, DCGAN, and DDPM, with the PIV dataset on the right for comparison.
(a) Mean local velocity fluctuations
(b) Variance of local velocity fluctuations
Figure 13: Comparison of (a) the mean local velocity fluctuation magnitude and (b) the variance of the local velocity fluctuations, evaluated over the wake region comprising 25% of the wake flow behind the array of seven cylinders in the 2V8H test case, extending to $12.6D$ within the numerical domain (see figure˜6(b)). Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
Figure 14: A Kiviat diagram comparing the three generative models for the LES flow around a cylinder dataset. Inference time and memory use are measured per sample, and the errors are measured by taking a mean square error between the respective error curves, and are then placed on a log-scale.
Figure 15: A Kiviat diagram comparing the three generative models for the PIV dataset 4V2H. Inference time and memory use are measured per sample, and the errors are measured by taking a mean square error between the respective error curves, and are then placed on a log-scale.